# Eigenvalue Gaps of the Laplacian of Random Graphs

### v1

@@ 5.1. Decomposition of vectors
This can be written as (here $x_{ij}$ are the entries of the adjacency matrix)
@@
where $f$ depends on the $a_{i}$ but not on the $x_{ij}$  [^5] This $f$ comes from the part ${\mathbf{E}}L_{n}$ in the definition of $\overline{L}_{n}$..
Note that there is also the part for ${\mathbf{v}}_{2}^{\prime}$ but we can fix the randomness over its vector support.
Hence for all $i\in I_{j_{0}}$, because the sum $x_{i1}+x_{i2}+\dots+x_{i|J_{1}|}$ takes value in [±$[\pm\Theta(\sqrt{\lambda_{n}n})]$ by Central Limit Theorem,
${\mathbb{P}}(((\overline{L}_{n}-\lambda_{0}){\mathbf{v}}^{\prime})_{i}\leq 2\sqrt{\log n}(\delta_{n}t_{0}^{8\lambda_{n}^{-1}})\leq 1/2,$
provided that $\delta_{n}^{\prime}\sqrt{\lambda_{n}}$ is sufficiently larger than $\sqrt{\log n}(\delta_{n}t_{0}^{\lambda_{n}^{-1}})$, which is satisfied by (30).
Taking over all $i\in\cup_{i}I_{j_{0}}$, by independence we hence obtain
${\mathbb{P}}\Big{(}((\overline{L}_{n}-\lambda_{0}){\mathbf{v}}^{\prime})_{j}\leq 2\sqrt{\log n}(\delta_{n}t_{0}^{8\lambda_{n}^{-1}}),i\in\cup_{i}I_{j_{0}}\Big{)}\leq(1/2)^{(n/2-\alpha_{n}n)}.$
Therefore, by tensorization (Lemma 3.3), we hence obtain
@@
Taking union bound over $D^{\alpha_{n}n}\times D^{\lambda_{n}^{-1}}\leq n^{Cn/(\log n)^{2}}$ choices of ${\mathbf{v}}^{\prime}$, we see that
@@
It thus follows that, as ${\mathbf{v}}$ is almost orthogonal to $\mathbf{1}$ because ${\mathbf{v}}$ is chosen as a non-trivial eigenvector (that is $\|{\mathbf{v}}-a\mathbf{1}/\sqrt{n}\|$ has order 1), $\|{\mathbf{v}}_{2}^{\prime}-a\mathbf{1}/\sqrt{n}\|_{2}$ is of order 1. By triangle inequality,
$\|(\overline{L}_{n}-\lambda_{0})({\mathbf{v}}^{\prime}-a\mathbf{1}/\sqrt{n}\|_{2}=O(\sqrt{\log n}(\delta_{n}t_{0}^{\lambda_{n}^{-1}})\sqrt{n})+\sqrt{n}\log n(\delta_{n}^{\prime})\ll\sqrt{n}/\log n,$
@@
On the other hand, by the Levy anti-concentration bound (see for instance [RV08, Lemma 2.6]), there exists $p<1$ such that
$\rho_{1/2}(\sum_{j\in\operatorname{supp}({\mathbf{v}}_{2}^{\prime})}x_{ij}(v_{j}^{\prime}-a/\sqrt{n}))<p.$
In particularly, by tensorization,
@@
taking union bound over the choices of ${\mathbf{v}}^{\prime}$, we obtain $D^{\lambda_{n}^{-1}}D^{\alpha_{n}n}p^{n-2\alpha_{n}n}$, which is small as in Subcase 1.1.
We now move to the harder scenario.
@@
Our treatment here is guided by the discussion from Subsection 2.5. We first pass to the approximated vector ${\mathbf{v}}^{\prime}$ and consider the event (29). Assume that $I_{s}$ consists of $i_{0}\leq\lambda_{n}^{-1}$ segments. We now describe a function $f$ from subsets of $[n]$ (that can occur as $I_{s}$) to subsets of $[n]$. We describe $F=f(I_{s})$, and write $I_{m}$ for $[n]\setminus I_{s}$, but note that $F$ does not depend on ${\mathbf{v}}$.
@@
In either case, we have $|F|=|I_{m}|$ and if $I_{m}$ is non-empty, it contains some $J_{i}$ that does not intersect $F$. Therefore, we create a square submatrix as well as an i.i.d. thin submatrix.
Let $c^{\ast}_{n}$ be chosen as
$c^{\ast}_{n}=\log^{-2}n$ (31)
Let ${\mathcal{E}}_{drop}$ be the event that there is a square submatrix $M_{A\times B}$ of $M=L_{n}$ of dimension $|A|=|B|$ at least $\lfloor n^{1-\delta_{0}}\rfloor$ for a sufficiently small constant $\delta_{0}$ so that $n^{1-\delta_{0}}<\lambda_{n}n$. with the $c^{\ast}_{n}|A|$-least singular value at least $c_{0}c^{\ast}_{n}\sqrt{|A|}$. Theorem 2.6 tells us that ${\mathbb{P}}({\mathcal{E}}_{drop})\leq e^{O(n)}e^{-\Theta(\lfloor n^{1-\delta_{0}}\rfloor)\sqrt{n}}$ (more specifically $e^{c^{\ast}_{n}n\log n}e^{-\Theta(c^{\ast}_{n}\lfloor n^{1-\delta_{0}}\rfloor)\sqrt{n}}$) and thus ${\mathbb{P}}({\mathcal{E}}_{drop})\leq e^{-c_{drop}n^{3/2-2\delta_{0}}},$ for some $c_{drop}>0$.
We wish to bound the probability of $\mathcal{E}_{2}\setminus{\mathcal{E}}_{drop}$. Since for ${\mathbf{v}}$ causing $\mathcal{E}_{2}$, we have $|I_{m}|>\lambda_{n}n$.
@@
Given subsets $S$ (for $I_{m}$), and $I_{m}^{\prime}\subset I_{m}$ of size $|I_{m}^{\prime}|=|I_{m}|-c^{\ast}_{n}|I_{m}|$, let $\mathcal{E}_{2.1}(S,{{{{I^{\prime}_{m}}}}}))$ be the event there exists ${\mathbf{v}}^{\prime}$ such that the above holds, and $\|(\overline{L}_{n}-\lambda_{0}){\mathbf{v}}^{\prime}\|_{2}$ is $O(\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|})$.
@@
For all $S,{{{{I^{\prime}_{m}}}}}$ such that $\mathcal{E}_{2.1}(S,{{{{I^{\prime}_{m}}}}})$ is defined, we have
@@
Proof.
(of Lemma 5.4) We fix a choice of ${\mathbf{v}}_{s}^{\prime}$ and ${\mathbf{v}}_{I_{m}\setminus{{I^{\prime}_{m}}}}^{\prime}$. In what follows
@@
Recall that the matrix $M_{12}$ is near isometry. Let $H_{12}$ be an $(|I_{m}|-c^{\ast}_{n}|I_{m}|)\times|I_{m}|$ matrix where $H_{12}M_{12}|_{I_{m}^{\prime}}=I_{I_{m}^{\prime}}$ (i.e. left inverse). By above (i.e. via. the application of Theorem 2.6), we have
@@
Now we can rewrite $\|\overline{L}_{n}{\mathbf{v}}^{\prime}\|_{2}\leq\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|}$ as $\|M_{1}^{\prime}{\mathbf{v}}_{s}^{\prime\prime}+M_{2}^{\prime}{\mathbf{v}}_{m}^{\prime\prime}\|_{2}\leq\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|}$, where
${\mathbf{v}}_{s}^{\prime\prime}=({\mathbf{v}}_{s}^{\prime},{\mathbf{v}}_{I_{m}\backslash I_{m}^{\prime}}^{\prime});{\mathbf{v}}_{m}^{\prime\prime}={\mathbf{v}}_{I_{m}^{\prime}}^{\prime}$
as
$\|(M_{11}{\mathbf{v}}_{s}^{\prime\prime},M_{12}{\mathbf{v}}_{s}^{\prime\prime})+(M_{21}{\mathbf{v}}_{m}^{\prime\prime},M_{22}{\mathbf{v}}_{m}^{\prime\prime})\|_{2}\leq\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|}.$ (32)
In particularly, $\|M_{11}{\mathbf{v}}_{s}^{\prime\prime}+M_{12}{\mathbf{v}}_{m}^{\prime\prime}\|_{2}\leq\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|}$, and after applying $H_{12}$ we have that
$\|H_{12}M_{11}{\mathbf{v}}_{s}^{\prime\prime}+{\mathbf{v}}_{m}^{\prime\prime}\|_{2}\leq\sqrt{n}\log n\times r_{n}c_{0}^{-1}(c^{\ast}_{n})^{-1}\sqrt{|I_{s}|}/\sqrt{|I_{m}|}\leq\sqrt{n}\log n\times\lambda_{n}^{-1/2}r_{n}c_{0}^{-1}(c^{\ast}_{n})^{-1}(\frac{1-c_{m}}{c_{m}})^{1/2}$
@@
where $|I_{m}|=c_{m}n$, with $\lambda_{n}/8<c_{m}<1-\lambda_{n}$. Here we note that $\delta_{n}$ is sufficiently small given all other parameters, and so the RHS bound will be small.
In other words, ${\mathbf{v}}_{m}^{\prime\prime}$ is almost determined if we fix $M_{11},M_{12}$ and ${\mathbf{v}}_{s}^{\prime\prime}$. As of this point, we will replace ${\mathbf{v}}_{m}^{\prime\prime}$ by ${\mathbf{u}}_{m}^{\prime\prime}=-H_{12}M_{11}{\mathbf{v}}_{s}^{\prime\prime}$. The latter vector is fixed if we condition on $M_{11},M_{12}$ and ${\mathbf{v}}_{s}^{\prime\prime}$. Note that
@@
Note that $M_{21}$ is the matrix $L_{[n]\backslash F\times(I_{s}\cup(I_{m}\backslash I_{m}^{\prime}))}$ and $M_{22}$ is the matrix $L_{[n]\backslash F\times I_{m}^{\prime}}$. We let $J_{*}$ be one of the $J_{i}$ that is in $I_{m}^{\prime}$ but has no intersection with $F$, where we will perform small ball probability and tensorization.
@@
The $x_{ij}$ for $i\in([n]\setminus F)\setminus J_{*}$ and $j\in J_{*}$ are all independent. We further condition on $x_{ij}$ for $i\in([n]\setminus F)\setminus J_{*}$ and $j\not\in J_{*}$. Since $J_{*}\cap F=\emptyset$, none of the $x_{ij}$ for $i\in([N]\setminus F)\setminus J_{*}$ and $j\in J_{*}$ have been conditioned on. Thus after our conditioning, the probability of (33) holding is at most $\rho_{L,\log n\times K_{n}r_{n}}({\mathbf{v}}_{J_{*}}^{\prime}).$ We have $n-|F|-|J_{*}|$ independent such inequality. Thus the probability that, for a given ${\mathbf{v}}_{s}$ and ${\mathbf{v}}_{I_{m}\setminus{{I^{\prime}_{m}}}},$ we have the event in the claim, is at most (by Lemma 3.3)
$\rho_{L,\log n\times K_{n}r_{n}}({\mathbf{v}}_{J_{*}}^{\prime})^{n-|F|-|J_{*}|}.$
Since $J_{*}$ is one of the $J_{i}$ that is a subset of $I_{m}$, by definition of $I_{m}$ and the highest multiplicity of coordinates of ${\mathbf{v}}_{J_{*}}$ is $<|J_{*}|-\alpha_{n}\lambda_{n}n$. More precisely, by the definition of non-sparse interval (25), for any $a$ there are at least $\alpha_{n}\lambda_{n}n$ entries $i\in J_{*}$ such that (noting that over the mix part, $v_{i}=v_{i}^{\prime}$)
@@
Theorem 5.1 (more specifically (21), and also in the spirit of Remark 5.2) imply that, with $r=t_{0}^{1/2}r_{n}$ and $R=\sqrt{\log n}K_{n}r_{n}$
$\rho_{L,\log n\times K_{n}r_{n}}({\mathbf{v}}_{J_{*}}^{\prime})=O(\frac{\log n\times K_{n}}{t_{0}^{1/2}}\frac{1}{\sqrt{\alpha_{n}\lambda_{n}n}})=O(\frac{\log^{3/2}n\times\alpha_{n}^{-1/2}\lambda_{n}^{-3/2}c_{0}^{-1}(c^{\ast}_{n})^{-1}}{t_{0}^{1/2}})=O(\frac{1}{t_{0}^{1/4}}).$
Note that with the choices of parameters $\lambda_{n}$ from Theorem 2.3, $c^{\ast}_{n}$ from (31), $\alpha_{n}$ from (24) and $t_{0}$ from (23) we see the that above is much smaller than 1.
We have $|J_{*}|\leq\lambda_{n}n/2$, $|F|=|M|$. By tensorization (Lemma 3.3), the probability that, for a given ${\mathbf{v}}_{s}^{\prime}$ and ${\mathbf{v}}_{I_{m}\setminus{{I^{\prime}_{m}}}}^{\prime}$, we have the event in the claim is at most
@@
Hence we obtain the following upper bound for the event ${\mathcal{E}}_{2.1}$ (given the choice of $t_{0}$ from (23))
@@
The event ${\mathcal{E}}_{2}$ is the union of ${\mathcal{E}}_{drop}$ with $\mathcal{E}_{2.1}(S,{{I^{\prime}_{m}}},{{{{I^{\prime}_{s}}}}})$ over $S,{{I^{\prime}_{m}}},{{{{I^{\prime}_{s}}}}}$ with $|S|-|{{I^{\prime}_{m}}}|=c_{\ast}|I_{m}|$ and $n-|S|\geq\lambda_{n}n$ and $|S|$ a possible value of $I_{m}$. There are at most $2^{\lambda_{n}^{-1}}$ possible values of $I_{m}$ (and hence $S$). Taking union bound over all the choices, and over all $I$ (of Theorem 2.3) of size $\lambda_{n}n$ we obtain an entropy bound say at most $4^{n}$, with room to spare. We obtain the following bound for the event (29) within Case 2
@@
6. No-structure delocalization: proof of Theorem 2.4 for the approximated vectors
@@
We also let ${\mathcal{E}}_{iid-norm}$ be the event that the matrix $\overline{L}_{n}$ with zeros on the diagonal (instead of the entries of ${\operatorname{diag}}(D_{n}-{\mathbf{E}}D_{n})$) has norm $O(\sqrt{n})$. It is well known that ${\mathcal{E}}_{iid-norm}$ holds with probability $1-e^{{\Theta(n)}}$.
Throughout this section, if not specified otherwise, we will condition on these two overwhelming events ${\mathcal{E}}_{non-gap}$ and ${\mathcal{E}}_{iid-norm}$.
@@
Any subvectors ${\mathbf{v}}_{J}$ of ${\mathbf{v}}$ of size $|J|=n_{0}/\sqrt{\log\log n}$ satisfies
@@
Assume that there exists an approximated vector ${\mathbf{v}}$ and an index $I$ such that
@@
where $D_{0}\asymp n^{A}$, for which under ${\mathcal{E}}_{non-gap}$, by Theorem 2.3 and Lemma 3.5 we have that $A\geq 1/2$, for appropriate choices of $\kappa,\gamma$ (such as $\kappa=n^{1/3}$ and $\gamma=c_{n}^{3}$ – which satisfy the conditions of Lemma 3.5 with $c_{0}$ a sufficiently small constant, and with $c_{1}=1/(\log n)^{A_{0}/c_{0}}$ and say $C=10$ respectively).

### v2

@@ 5.1. Decomposition of vectors
This can be written as
@@
where $x_{ij}$ are the entries of the adjacency matrix, and $f$ depends on the $a_{i}$ but not on the $x_{ij}$ [^6] This $f$ comes from the part ${\mathbf{E}}L_{n}$ in the definition of $\overline{L}_{n}$..
Note that there is also the part for ${\mathbf{v}}_{2}^{\prime}$, but we can fix the randomness over its vector support.
Hence, for all $i\in I_{j_{0}}$, because the sum $x_{i1}+x_{i2}+\dots+x_{i|J_{1}|}$ lies in [±$[\pm\Theta(\sqrt{\lambda_{n}n})]$ by the Central Limit Theorem,
${\mathbb{P}}\left(((\overline{L}_{n}-\lambda_{0}){\mathbf{v}}^{\prime})_{i}\leq 2\sqrt{\log n}\left(\delta_{n}t_{0}^{8\lambda_{n}^{-1}}\right)\right)\leq 1/2,$
provided that $\delta_{n}^{\prime}\sqrt{\lambda_{n}}$ is sufficiently larger than $\sqrt{\log n}(\delta_{n}t_{0}^{\lambda_{n}^{-1}})$; this condition is satisfied by (31).
Combining the above conclusion for all $i\in\cup_{j_{0}}I_{j_{0}}$ yields
${\mathbb{P}}\Big{(}((\overline{L}_{n}-\lambda_{0}){\mathbf{v}}^{\prime})_{j}\leq 2\sqrt{\log n}(\delta_{n}t_{0}^{8\lambda_{n}^{-1}}),i\in\cup_{j_{0}}I_{j_{0}}\Big{)}\leq(1/2)^{(n/2-\alpha_{n}n)}.$
Therefore, by tensorization (Lemma 3.3), we obtain
@@
Taking union bound over $D^{\alpha_{n}n}\times D^{8\lambda_{n}^{-1}}\leq n^{Cn/(\log n)^{2}}$ choices of ${\mathbf{v}}^{\prime}$, we see that
@@
Recall that ${\mathbf{v}}$ is almost orthogonal to $\mathbf{1}$, since ${\mathbf{v}}$ is chosen as an non-trivial approximate eigenvector. Therefore, $\|{\mathbf{v}}-a\mathbf{1}/\sqrt{n}\|$ has order 1, so that $\|{\mathbf{v}}_{2}^{\prime}-a\mathbf{1}/\sqrt{n})\|_{2}$ is of order 1. By triangle inequality,
$\|(\overline{L}_{n}-\lambda_{0})({\mathbf{v}}^{\prime}-a\mathbf{1}/\sqrt{n})\|_{2}=O(\sqrt{\log n}(\delta_{n}t_{0}^{\lambda_{n}^{-1}})\sqrt{n})+\sqrt{n}\log n(\delta_{n}^{\prime})\ll\sqrt{n}/\log n,$
@@
On the other hand, by the Levy anti-concentration bound (see for instance [58, Lemma 2.6]), there exists a constant $p<1$ such that
$\rho_{1/2}\left(\sum_{j\in\operatorname{supp}({\mathbf{v}}_{2}^{\prime})}x_{ij}(v_{j}^{\prime}-a/\sqrt{n})\right)<p.$
In particular, by tensorization,
@@
Taking a union bound over the choices of ${\mathbf{v}}^{\prime}$, we obtain $D^{\lambda_{n}^{-1}}D^{\alpha_{n}n}p^{n-2\alpha_{n}n}$, which is small as in Subcase 1.1.
Finally, taking union bound over $\binom{n}{\lfloor\lambda_{n}n\rfloor}$ choices for $I$, we see that, with $\lambda_{n}$ from (22), the union bounds from both Subcase 1.1 and Subcase 1.2 are of order $e^{-\Theta(n)}$.
We now move to Case 2, which is more involved.
@@
Our treatment here is guided by the discussion from Subsection 2.5. We first pass to the approximate eigenvector ${\mathbf{v}}^{\prime}$ and consider the event (30). Assume that $I_{s}$ consists of $i_{0}\leq 8\lambda_{n}^{-1}$ segments. We now describe a function $f$ from subsets of $[n]$ (that can occur as $I_{s}$) to subsets of $[n]$. We describe $F=f(I_{s})$, and write $I_{m}$ for $[n]\setminus I_{s}$, but note that $F$ does not depend on ${\mathbf{v}}$.
@@
In either case, we have $|F|=|I_{m}|$ and if $I_{m}$ is non-empty, it contains some $J_{i}$ that does not intersect $F$. Therefore, we create a square submatrix as well as an i.i.d. thin submatrix out of these indices.
Let
$c^{\ast}_{n}=\log^{-2}n,$ (32)
and let ${\mathcal{E}}_{drop}$ be the event that there is a square submatrix $M_{A\times B}$ of $M=L_{n}$ of dimension $|A|=|B|\geq\lfloor n^{1-\delta_{0}}\rfloor$ for a sufficiently small constant $\delta_{0}$ so that $n^{1-\delta_{0}}<\lambda_{n}n$, with the $c^{\ast}_{n}|A|$-least singular value at least $c_{0}c^{\ast}_{n}\sqrt{|A|}$. Theorem 2.6 tells us that ${\mathbb{P}}({\mathcal{E}}_{drop})\leq e^{O(n)}e^{-\Theta(\lfloor n^{1-\delta_{0}}\rfloor)\sqrt{n}}$ (more specifically, $e^{c^{\ast}_{n}n\log n}e^{-\Theta(c^{\ast}_{n}\lfloor n^{1-\delta_{0}}\rfloor)\sqrt{n}}$), and thus ${\mathbb{P}}({\mathcal{E}}_{drop})\leq e^{-c_{drop}n^{3/2-2\delta_{0}}},$ for some $c_{drop}>0$.
We wish to bound the probability of $\mathcal{E}_{2}\setminus{\mathcal{E}}_{drop}$. Recall that for ${\mathbf{v}}$ causing $\mathcal{E}_{2}$, we have $|I_{m}|>\lambda_{n}n/8$.
@@
Given subsets $S$ (for $I_{m}$), and $I_{m}^{\prime}\subset I_{m}$ of size $|I_{m}^{\prime}|=|I_{m}|-c^{\ast}_{n}|I_{m}|$, let $\mathcal{E}_{2.1}(S,{{{{I^{\prime}_{m}}}}}))$ be the event that there exists ${\mathbf{v}}^{\prime}$ such that the above holds, and $\|(\overline{L}_{n}-\lambda_{0}){\mathbf{v}}^{\prime}\|_{2}$ is $O(\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|})$.
@@
For all $S$ and $I_{m}^{\prime}$ such that $\mathcal{E}_{2.1}(S,I^{\prime}_{m})$ is defined, we have that
@@
Proof of Lemma 5.4.
We fix a choice of ${\mathbf{v}}_{s}^{\prime}$ and ${\mathbf{v}}_{I_{m}\setminus{{I^{\prime}_{m}}}}^{\prime}$. In what follows, we denote $M_{11},M_{12},M_{21},M_{22}$ as follows:
@@
Recall that the matrix $M_{12}$ is near isometry. Let $H_{12}$ be an $(|I_{m}|-c^{\ast}_{n}|I_{m}|)\times|I_{m}|$ matrix where $H_{12}M_{12}|_{I_{m}^{\prime}}=I_{I_{m}^{\prime}}$ (i.e. $H_{12}$ is a left inverse of $M_{12}$). By the above (i.e. via the application of Theorem 2.6), we have
@@
If we write
${\mathbf{v}}_{s}^{\prime\prime}=({\mathbf{v}}_{s}^{\prime},{\mathbf{v}}_{I_{m}\backslash I_{m}^{\prime}}^{\prime})\quad\text{and}\quad{\mathbf{v}}_{m}^{\prime\prime}={\mathbf{v}}_{I_{m}^{\prime}}^{\prime},$
so that
$\|\overline{L}_{n}{\mathbf{v}}^{\prime}\|_{2}=\|(M_{11}{\mathbf{v}}_{s}^{\prime\prime},M_{12}{\mathbf{v}}_{s}^{\prime\prime})+(M_{21}{\mathbf{v}}_{m}^{\prime\prime},M_{22}{\mathbf{v}}_{m}^{\prime\prime})\|_{2}\leq\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|}.$ (33)
Then we have that $\|M_{1}^{\prime}{\mathbf{v}}_{s}^{\prime\prime}+M_{2}^{\prime}{\mathbf{v}}_{m}^{\prime\prime}\|_{2}\leq\sqrt{n}\log n\times r_{n}\sqrt{|I_{s}|}$. In particular, after applying $H_{12}$ we have that
$\|H_{12}M_{11}{\mathbf{v}}_{s}^{\prime\prime}+{\mathbf{v}}_{m}^{\prime\prime}\|_{2}\leq\sqrt{n}\log n\times r_{n}c_{0}^{-1}(c^{\ast}_{n})^{-1}\sqrt{|I_{s}|}/\sqrt{|I_{m}|}\leq\sqrt{n}\log n\times\lambda_{n}^{-1/2}r_{n}c_{0}^{-1}(c^{\ast}_{n})^{-1}\left(\frac{1-c_{m}}{c_{m}}\right)^{1/2}$
@@
where $|I_{m}|=c_{m}n$, with $c_{m}$ satisfying $\lambda_{n}/8<c_{m}<1-\lambda_{n}$. Here we note that $\delta_{n}$ is sufficiently small given all other parameters, and so the RHS bound will be small.
In other words, ${\mathbf{v}}_{m}^{\prime\prime}$ is almost determined if we fix $M_{11},M_{12}$ and ${\mathbf{v}}_{s}^{\prime\prime}$. At this point, we will replace ${\mathbf{v}}_{m}^{\prime\prime}$ by ${\mathbf{u}}_{m}^{\prime\prime}=-H_{12}M_{11}{\mathbf{v}}_{s}^{\prime\prime}$. The latter vector is fixed if we condition on $M_{11},M_{12}$, and ${\mathbf{v}}_{s}^{\prime\prime}$. Note that
@@
Note that $M_{21}$ is the matrix $L_{[n]\backslash F\times(I_{s}\cup(I_{m}\backslash I_{m}^{\prime}))}$ and $M_{22}$ is the matrix $L_{[n]\backslash F\times I_{m}^{\prime}}$. We let $J_{*}$ be one of the $J_{i}$ that is in $I_{m}^{\prime}$ but has no intersection with $F$. It is over these coordinates that we will calculate the small ball probabilities and subsequently apply tensorization.
@@
Notice that the collection $\left\{x_{ij}\colon i\in([n]\setminus F)\setminus J_{*},j\in J_{*}\right\}$ is independent. We further condition on $x_{ij}$ for $i\in([n]\setminus F)\setminus J_{*}$ and $j\not\in J_{*}$. Since $J_{*}\cap F=\emptyset$, none of the $x_{ij}$ for $i\in([N]\setminus F)\setminus J_{*}$ and $j\in J_{*}$ have been conditioned on. Thus, after our conditioning, the probability of (34) holding is at most $\rho_{L,\log n\times K_{n}r_{n}}({\mathbf{v}}_{J_{*}}^{\prime}).$ Thus, by tensorizing (Lemma 3.3) over the $n-|F|-|J^{*}|$ independent rows, the probability that we have the event in the claim for a given ${\mathbf{v}}_{s}$ and ${\mathbf{v}}_{I_{m}\setminus{{I^{\prime}_{m}}}}$ is at most
$\rho_{L,\log n\times K_{n}r_{n}}({\mathbf{v}}_{J_{*}}^{\prime})^{n-|F|-|J_{*}|}.$ (35)
Since $J_{*}$ is one of the $J_{i}$ that is a subset of $I_{m}$. Therefore, by the definition of non-sparse interval (26), for any $a$ there are at least $\alpha_{n}\lambda_{n}n$ coordinates $i\in J_{*}$ such that (noting that over the mix part, $v_{i}=v_{i}^{\prime}$)
@@
Theorem 5.1 (more specifically (21), and also in the spirit of Remark 5.2) implies that, with $r=t_{0}^{1/2}r_{n}$ and $R=\sqrt{\log n}K_{n}r_{n}$,
$\rho_{L,\log n\times K_{n}r_{n}}({\mathbf{v}}_{J_{*}}^{\prime})=O\left(\frac{\log n\times K_{n}}{t_{0}^{1/2}}\frac{1}{\sqrt{\alpha_{n}\lambda_{n}n}}\right)=O\left(\frac{\log^{3/2}n\times\alpha_{n}^{-1/2}\lambda_{n}^{-3/2}c_{0}^{-1}(c^{\ast}_{n})^{-1}}{t_{0}^{1/2}}\right)=O\left(\frac{1}{t_{0}^{1/4}}\right).$
Note that with the choices of parameters $\lambda_{n}$ from Theorem 2.3, $c^{\ast}_{n}$ from (32), $\alpha_{n}$ from (25), and $t_{0}$ from (24), we see the that above is much smaller than 1. We have $|J_{*}|\leq\lambda_{n}n/2$, $|F|=|I_{m}|$, therefore combining the above with (35), we have that the probability of the event in the claim for a given ${\mathbf{v}}_{s}$ and ${\mathbf{v}}_{I_{m}\setminus{{I^{\prime}_{m}}}}$ is at most
@@
Hence we obtain the following upper bound for the event ${\mathcal{E}}_{2.1}$ (given the choice of $t_{0}$ from (24)):
@@
The event ${\mathcal{E}}_{2}$ is the union of ${\mathcal{E}}_{drop}$ with $\mathcal{E}_{2.1}(S,{{I^{\prime}_{m}}},{{{{I^{\prime}_{s}}}}})$ over $S,{{I^{\prime}_{m}}},{{{{I^{\prime}_{s}}}}}$, with $|S|-|{{I^{\prime}_{m}}}|=c_{\ast}|I_{m}|$, and $n-|S|\geq\lambda_{n}n$ and $|S|$ a possible value of $I_{m}$. There are at most $2^{8\lambda_{n}^{-1}}$ possible values of $I_{m}$ (and hence $S$). Taking a union bound over all the choices, and over all $I$ (of Theorem 2.3) of size $\lambda_{n}n$ we obtain an entropy bound at most $4^{n}$. We obtain the following bound for the event (30) within Case 2
@@
6. No-Structure Delocalization: Proof of Theorem 2.4 for the Approximate Eigenvectors
@@
We also let ${\mathcal{E}}_{i.i.d.-norm}$ be the event that the matrix $\overline{L}_{n}$ with zeros on the diagonal (instead of the entries of ${\operatorname{diag}}(D_{n}-{\mathbf{E}}D_{n})$) has norm $O(\sqrt{n})$. It is well known that ${\mathcal{E}}_{i.i.d.-norm}$ holds with probability $1-e^{{\Theta(n)}}$.
Throughout this section, if not specified otherwise, we will condition on these two overwhelming events ${\mathcal{E}}_{non-gap}$ and ${\mathcal{E}}_{i.i.d.-norm}$.
@@
Any subvector ${\mathbf{v}}_{J}$ of ${\mathbf{v}}$ of size $|J|=n_{0}/\sqrt{\log\log n}$ satisfies
@@
Assume that there exists an approximate eigenvector ${\mathbf{v}}$ and an index $I$ such that
@@
where $D_{0}\asymp n^{A}$. Under ${\mathcal{E}}_{non-gap}$, by Theorem 2.3 and Lemma 3.5, we have that $A\geq 1/2$, for appropriate choices of $\kappa,\gamma$ (such as $\kappa=n^{1/3}$ and $\gamma=c_{n}^{3}$ – which satisfy the conditions of Lemma 3.5 with $c_{0}$ a sufficiently small constant, and with $c_{1}=1/(\log n)^{A_{0}/c_{0}}$ and $C=10$ respectively).